OpenStudy (anonymous):

Solve (5x – 8) + 2 = 5x – (8 + 2). Answer x = 6/5 x = 0 This is an identity and is true for all real numbers. There is no solution.

OpenStudy (anonymous):

@iNeedTheHelp

OpenStudy (anonymous):

ok the answer is no solution

OpenStudy (anonymous):

Yup! dpgreen is right! When it comes out, it comes out as 0. :)

OpenStudy (anonymous):

5x-8+2=5x-8-2 subtract 5x from the left -8+2=-8-2 -6 does not equal 6